Complete combustion of 4.00 g of a hydrocarbon produced 12.3 g of CO2 and 5.68 g of H2O. What is the empirical formula for the hydrocarbon? Insert subscripts as necessary.
Answer #1

Complete combustion of 4.00 g of a hydrocarbon produced 12.3 g of CO2 and 5.68 g of H2O. What is the empirical formula for the hydrocarbon?

moles of CO2 produced : 12.3 g / 44 g/mol = 0.28 moles

moles of H2O produced : 5.68 g/ 18 g/mol = 0.316 moles

since ,

CxHy + (x+ y/4) = x CO2 + y/2 H2O

thus, x = 0.28 and y = 0.63

on removing fractions, ===> x = 4 , H = 9

empirical formula : C4H9

( empirical formula of a chemical compound is the simplest positive integer ratio of atoms present in a compound ).